# Inter-Speaker Relative Cues for Text-Guided Target Speech Extraction

### v1

@@ Abstract
We propose a novel approach that utilize inter-speaker rela- tive cues for distinguishing target speakers and extracting their voices from mixtures. Continuous cues (e.g., temporal order, age, pitch level) are grouped by relative differences, while dis- crete cues (e.g., language, gender, emotion) retain their cate- gories. Relative cues offers greater flexibility than fixed speech attribute classification, facilitating much easier expansion of text-guided target speech extraction datasets. Our experiments show that combining all relative cues yields better performance than random subsets, with gender and temporal order being the most robust across languages and reverberant conditions. Ad- ditional cues like pitch level, loudness, distance, speaking dura- tion, language, and pitch range also demonstrate notable benefit in complex scenarios. Fine-tuning pre-trained WavLM Base+ CNN encoders improves overall performance over the baseline of using only a Conv1d encoder.
@@ 1 Introduction
Compared to the aforementioned cues, text is more accessible, making it a practical and flexible option. With advances in language models, several studies have used natural language descriptions to extract specific sound events or musical instruments from audio mixtures, achieving impressive results [6, 7, 8]. Building on this, other research has employed text descriptions to isolate a specific speaker’s voice from a speech mixture. LLM-TSE [9] was the first to use natural language descriptions powered by a large language model to extract semantic information from text queries (such as a description of single speech attribute, e.g., gender, language) and integrate it into a speech separation network, demonstrating the practicality of text-guided TSE.
However, its focus on a single speech attribute limits its broader application. To overcome this, Huo et al. introduced a more flexible text-guided TSE model and developed the TextrolMix dataset, which provides rich descriptions of speech attributes [10]. This dataset offers two types of contextual clues: (1) a natural language description of the target speech’s speaking style and (2) a reference audio that shares specific style attributes with the target speech, accompanied by a corresponding text prompt. Each utterance is annotated with six attributes: speaker identity, emotion, pitch level, gender, accent, and tempo (speaking rate). The speaker identity is represented by an enrollment speech sample from the target speaker, aligned with the relevant text prompt.
@@
To address these limitations, we incorporate a broad range of speech attributes and create a two-speaker mixture dataset, including language, gender, transcription, emotion, temporal order, age, speaking rate, duration, pitch level (mean $F_{0}$), pitch range ($F_{0}$ span), loudness, and speaker-to-microphone distance. The contributions of our work are as follows: (i) We introduce inter-speaker relative cues, which categorize the collected attributes based on the differences between the target and interference speakers’ voices. The relative cues provide a way to overcome the fixed categorization of speech attributes and make the dataset much easier to expand. (ii) We conduct baseline experiment to evaluate their impact and contribution in distinguishing target speakers and extracting their voices. (iii) Since the addition of more cues can complicate the learning process, we propose utilizing pre-trained weights of WavLM Base+ CNN encoders for improved extraction performance. To facilitate future research in this direction, code for creating our dataset is available at: https://github.com/daiwangsnr.
@@ 2.2 Audio part
(iv) We simulate reverberant speech using the generated clean speech samples and gpuRIR, which models speakers in a room with reverberation effects. The room dimensions range from 9–11 m in length and width, 2.6–3.5 m in height, and reverberation times of 0.3–0.6 seconds. The microphone is positioned at the center of the room, while speaker-to-microphone distances vary from 0.3 to 1.5 m horizontally, with speaker heights between 1.6 and 1.9 m. We generate 10,000 RIR pairs for training, 1,000 for validation, and 1,000 for testing. For each mixing iteration, we randomly select RIRs ($rir_{1}$, $rir_{2}$) and convolve them with speech signals ($S_{1}^{\prime}$, $S_{2}^{\prime}$). The speech-to-interference ratio (SIR) between the reverberant speech samples is uniformly sampled from 0 to 6 dB before merging them into a single mixture.
@@ 2.3 Text Prompt Generation
The goal is to create generic text prompts that capture inter-speaker relative cues for target speech extraction. We start by compiling template prompts based on one or more relative cues, following a similar approach to [8]. Since individual relative cues are intended to specify the target speaker, cues classified as “same” or “similar” are excluded. The template prompt is generated by inserting actions and cues into structures like "Please <verb> <str(cues)>." or "Can you <verb> str(cues)?". Here, the verb represents an action, randomly selected from “extract”, “isolate”, and “separate”, while the str function converts a cue vector into a human-readable string by inserting a noun phrase, such as “speaker characterized by”. For instance, the cue vector (female, higher pitch level, faster speaking rate) can be expressed as: “the female speaker characterized by a higher pitch level and a faster speaking rate”
For each mixed speech sample in the TSE training, validation, and test sets, we generate multiple template prompts, including twelve individual cues, a randomly selected subset of cues (random cues), and a combination of all cues. To enhance template flexibility, we rephrase each prompt into five variations using ChatGPT-4o-mini, leveraging its efficiency and accuracy. We instruct the model to focus on action verbs such as “extract”, “separate” and “isolate” while avoiding terms like “identify” and “locate”. Additionally, we encourage the model to explore different sentence structures and synonyms for speech- or speaker-related keywords, such as “female”, “male”, “pitch range”, “speaking speed”, “quieter loudness”, “louder loudness”, “longer speaking duration”, “shorter speaking duration” and so on. Furthermore, we instruct the model to ensure the language remains simple and suitable for everyday use. After generating these generic prompts, we found that only a few outputs were incomplete or irrelevant. These were manually removed and regenerated.
@@ 3.2 Training and Network Configurations
The LLaMA 3.2 1B instruction model was fine-tuned using LoRA (rank 16, scaling weight 16, dropout 0.05), applied to query and key projection layers in self-attention. Training was performed using the negative SI-SDR loss with float16 mixed-precision and a batch size of 24. The AdamW optimizer was used with a learning rate of 1e-4, halved if validation loss did not decrease for three consecutive epochs. A linear warm-up was applied during the first 1,000 updates, and gradient normalization (with a maximum norm of 30) was employed to prevent gradient explosion. Models were trained with early stopping, with up to 100 epochs. Following [8], we used all five generic text prompts with a randomly selected subset of cues (random cues) for each mixture during training. For validation, only the fifth set of generic prompts with random cues were used.
@@ 3.3 Results and Discussion
We evaluate target speech extraction performance using SI-SDR improvement (SI-SDRi) and perceptual evaluation of speech quality (PESQ) [30]. The fifth set of generic prompts, including random, individual, and all cues per test sample, is used for evaluation. Table 3 presents the results when both models are trained with random cues.
@@
The results show the potential of training with only random cues, while using all cues leads to better performance by utilizing additional information during the inference stage. Notably, certain cues, such as gender and temporal order, exhibit strong discriminative power across different speech signals. These cues are language- and reverberation-independent and are more perceptible to humans than other individual cues. Cues like speaking duration, pitch level, and loudness also perform well across languages and in reverberant environments. The pitch range cue doesn’t work as well as pitch level, likely because reverberation—especially under strong conditions—distorts the boundaries of pitch changes. In contrast, the age category performs poorly, likely due to limited training samples and its sensitivity to both language and reverberation effects. Additionally, some age-annotated speech includes emotional speech, further complicating differentiation. Meanwhile, transcription, emotion, and speaking rate cues outperform age category, as they are primarily influenced by language variations. The comparison between the baseline and proposed method indicates that integrating pre-trained WavLM Base+ features improves overall SI-SDR performance, especially for language, transcription, gender and distance cues, while also enhancing speech perception quality across all conditions.
@@ 4 Conclusions and Future Work
This study explored rich relative cues between speakers for text-guided target speech extraction. Experimental Results suggest the effectiveness of incorporating multiple cues and using the pre-trained representation of WavLM Base+ CNN encoders. The findings further reveal that different cues have varying impacts and contributions in extracting target speakers’ voices. Among the cues, gender and temporal order proved the most robust in cross-lingual and reverberant conditions, while cues like pitch level, loudness, distance, speaking duration, language, and pitch range also contributed, yielding over or nearly 3 dB SI-SDR gains for both models. Age category underperformed due to limited data and sensitivity to complex environment. Future work will expand age-labeled data, refine the model or training strategy to boost overall performance, and assess generalization across different languages and reverberation conditions.

### v3

@@ Abstract
We propose a novel approach that utilizes inter-speaker relative cues to distinguish target speakers and extract their voices from mixtures. Continuous cues (e.g., temporal order, age, pitch level) are grouped by relative differences, while discrete cues (e.g., language, gender, emotion) retain their categorical distinctions. Compared to fixed speech attribute classification, inter-speaker relative cues offer greater flexibility, facilitating much easier expansion of text-guided target speech extraction datasets. Our experiments show that combining all relative cues yields better performance than random subsets, with gender and temporal order being the most robust across languages and reverberant conditions. Additional cues, such as pitch level, loudness, distance, speaking duration, language, and pitch range, also demonstrate notable benefits in complex scenarios. Fine-tuning pre-trained WavLM Base+ CNN encoders improves overall performance over the Conv1d baseline.
@@ 1 Introduction
Compared to the aforementioned cues, text is more accessible, making it a practical and flexible option. With advances in language models, several studies have used natural language descriptions to extract specific sound events or musical instruments from audio mixtures, achieving impressive results [6, 7, 8]. Building on this, recent study has employed text descriptions to isolate a specific speaker’s voice from a speech mixture. LLM-TSE [9] was the first to use natural language descriptions powered by a large language model to extract semantic information from text queries (such as a description of single speech attribute, e.g., gender, language) and integrate it into a speech separation network, demonstrating the effectiveness of text-guided TSE.
However, its focus on a single speech attribute limits its broader application. To overcome this, Huo et al. introduced a more flexible text-guided TSE model and developed the TextrolMix dataset, which provides rich descriptions of speech attributes [10]. This dataset offers two types of contextual clues: (1) a natural language description of the speaking style of the target speech and (2) a reference audio that shares specific style attributes with the target speech, accompanied by a corresponding text prompt. Each utterance is annotated with six attributes: speaker identity, emotion, pitch level, gender, accent, and tempo (speaking rate). The speaker identity is represented by an enrollment speech sample from the target speaker, aligned with the relevant text prompt.
@@
To address these limitations, we incorporate a broad range of speech attributes and create a two-speaker mixture dataset, including language, gender, transcription, emotion, temporal order, age, speaking rate, duration, pitch level (mean $F_{0}$), pitch range ($F_{0}$ span), loudness, and speaker-to-microphone distance. The contributions of our work are as follows: (i) We introduce inter-speaker relative cues, which categorize the collected attributes based on the differences between the target and interference speakers’ voices. The relative cues provide a way to overcome the fixed categorization of speech attributes and make the dataset much easier to expand. (ii) We conduct baseline experiment to evaluate the impact and contribution of these cues in extracting target speakers’ voices. (iii) Since the addition of more cues can complicate the learning process, we propose utilizing pre-trained weights of WavLM Base+ CNN encoders for improved extraction performance. To facilitate future research in this direction, code for creating our dataset is available at: https://github.com/daiwangsnr.
@@ 2.2 Audio part
(iv) We simulate reverberant speech using the generated clean speech samples and gpuRIR library [28], which models speakers in a room with reverberation effects. The room dimensions range from 9–11 m in length and width, 2.6–3.5 m in height, and reverberation times of 0.3–0.6 seconds. The microphone is positioned at the center of the room, while speaker-to-microphone distances vary from 0.3 to 1.5 m horizontally, with speaker heights between 1.6 and 1.9 m. We generate 10,000 RIR pairs for training, 1,000 for validation, and 1,000 for testing. For each mixing iteration, we randomly select RIRs ($rir_{1}$, $rir_{2}$) and convolve them with clean speech signals ($S_{1}^{\prime}$, $S_{2}^{\prime}$). The SIR between each reverberant speech pair is uniformly sampled from -6 to 6 dB before merging them into a single mixture.
@@ 2.3 Text Prompt Generation
The goal is to create generic text prompts that capture inter-speaker relative cues for target speech extraction. We start by compiling template prompts based on one or more relative cues, following a similar approach to [8]. Since individual relative cues are intended to specify the target speaker, cues classified as “same” or “similar” are excluded. The template prompt is generated by inserting actions and cues into structures like "Please <verb> <str(cues)>." or "Can you <verb> str(cues)?". Here, the verb represents an action, randomly selected from “extract”, “isolate”, and “separate”, while the str function converts a cue vector into a human-readable string by inserting a noun phrase, such as “speaker characterized by”. For instance, the cue vector (female, higher pitch level, faster speaking rate) can be expressed as: “the female speaker characterized by a higher pitch level and a faster speaking rate”.
For each mixed speech sample in the TSE training, validation, and test sets, we generate multiple template prompts, including twelve individual cues, a randomly selected subset of cues (random cues), and a combination of all individual cues (all cues). To enhance template flexibility, we rephrase each prompt into five variations using ChatGPT-4o-mini, leveraging its efficiency and accuracy. We instruct the model to focus on action verbs such as “extract”, “separate” and “isolate” while avoiding terms like “identify” and “locate”. In addition, we encourage the model to explore different sentence structures and synonyms for keywords that describe speech or speaker characteristics, such as “female”, “male”, “pitch range”, “speaking speed”, “quieter loudness”, “louder loudness”, “longer speaking duration”, “shorter speaking duration”, and so on. Furthermore, we instruct the model to ensure the language remains simple and suitable for everyday use. After generating these generic prompts, we found that only a few outputs were incomplete or irrelevant. These were manually removed and regenerated.
@@ 3.2 Training and Network Configurations
The LLaMA 3.2 1B instruction model was fine-tuned using LoRA (rank 16, scaling factor 16, dropout 0.05), applied to query and key projection layers in self-attention. Training was performed using the negative SI-SDR [31] loss with float16 mixed-precision and a batch size of 24. The AdamW optimizer was used with a learning rate of 1e-4, halved if validation loss did not decrease for three consecutive epochs. A linear warm-up was applied during the first 1,000 updates, and gradient normalization (with a maximum norm of 30) was employed to prevent gradient explosion. Models were trained with early stopping, with up to 100 epochs. Following [8], we used all five generic text prompts with a randomly selected subset of cues (random cues) for each mixture during training. For validation, only the fifth set of generic prompts with random cues were used.
@@ 3.3 Results and Discussion
We evaluate target speech extraction performance using SI-SDR improvement (SI-SDRi) and perceptual evaluation of speech quality (PESQ) [32]. The fifth set of generic prompts in the test set, which includes random, individual, and all cues for each test sample, is used for evaluation. Table 3 presents the results when both models are trained with random cues.
@@
The results show the potential of training with only random cues, while using all cues leads to better performance by utilizing additional information during the inference stage. Notably, certain cues, such as gender and temporal order, exhibit strong discriminative power across different speech signals. These cues are language- and reverberation-independent and are more perceptible to humans than other individual cues. Cues like speaking duration, pitch level, and loudness also perform well across languages and in reverberant environments. The pitch range cue doesn’t work as well as pitch level, likely because reverberation—especially under strong conditions—distorts the boundaries of pitch changes. In contrast, the age category performs poorly, likely due to limited training samples and its sensitivity to both language and reverberation effects. Additionally, some age-annotated speech samples exhibit emotion, further complicating differentiation. Meanwhile, transcription, emotion, and speaking rate cues outperform age category, as they are primarily influenced by language variations. The comparison between the baseline and proposed method indicates that integrating pre-trained WavLM Base+ features improves overall SI-SDRi, especially for language, transcription, gender, and distance cues, while also enhancing speech perception quality across all conditions.
@@ 4 Conclusions and Future Work
This study explored rich relative cues between speakers for text-guided target speech extraction. Experimental results suggest the effectiveness of incorporating multiple cues and leveraging pre-trained representations from WavLM Base+ CNN encoders. The findings also reveal that different cues have varying impacts and contributions in extracting target speakers’ voices. Among the cues, gender and temporal order proved the most robust in cross-lingual and reverberant conditions, while cues like pitch level, loudness, distance, speaking duration, language, and pitch range also contributed, yielding over or nearly 3 dB SI-SDR gains for both models. Age category underperformed due to limited data and sensitivity to complex environment. Future work will expand age-labeled data, refine the model or training strategy to boost overall performance, and assess generalization across different languages and reverberation conditions.